The number of quadratic equation which are unchanged by squaring their roots is

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

let x2 +bx +c=0 ............1

not let another equation whoose roots are square of this equation

so   X=x2   or   x=√X

 put value of x

  X +b√X  +c =0

 or X +c =-b√X

square

   X2 +c2 +2cX =b2X

   X2 +(2c-b2)X + c2 =0..................2

root of equation 2 is the square of root of equation 1

both equation will be same if their coefficient are in proportion

1/1  =b/(2c-b2)  =c/c 2

b= 2c-b2     ................3            

c=c2  ....................3

 from equation 3  c=0 or 1

 for c =0  b= 0 and -1

  for c=1  b= 1 and -2 

So, four combination are possible

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

All the best.
Regards,
Askiitians Experts
Badiuddin

Jay Lal
11 years ago

There are four such cases x=0 root 0

(x-1)2=0  root 1

x(x+1)=0  roots  0 and 1

x2+x+1=0 roots ω and ω2 

Panchin Ganeriwal
7 years ago
The four roots are:

Equations             roots
x2 = 0                       0,0
x (x-1)=0.                 1,0
(x-1)2 = 0.                1,1
x2+x+1=0              ω  ω2
I think u all must have found out the first three equarions , but the challenge was the fourth one